Toss a biased coin: for this coin the probability that it will show tail is 0.7. We toss coin twice. Find probability that we get no head
Answer by greenestamps(13198)



P(no head) = P(tails twice) = (0.7)(0.7) = 0.49

P(exactly one head) = P(tails then heads OR heads then tails) = (0.7)(0.3)+(0.3)(0.7) = .21+.21 = .42

P(two heads) = (0.3)(0.3) = 0.09

Note that the probabilities of all possible cases sum to 1, as they must:
0.49+0.42+0.09 = 1.00